What is the volume occupied by 28.0 g of oxygen, O2 at STP?


1
Expert's answer
2020-11-04T14:12:00-0500

The volume is


"V=nV_M,"

where


"n=\frac{m}{\Mu}."

Thus:


"V=\frac{m}{\Mu} V_M=\frac{28}{16+16}22.4=28.5\text{ L}."